Suppose a small economy has consumption spending of $6,500 when its income is $10,000. When income goes up to $11,000 then consumption spending is $7,250. What is the government spending multiplier in this economy? (Hint: Find the MPC first, then calculate the multiplier.) 2.85 4.00 1.53 7.00
Answer #1

Answer : Based on given information,

Changes in consumption spending = 7250 - 6500 = $750

Changes in income = 11000 - 10000 = $1,000

MPC = Changes in consumption spending / Changes in income

=> MPC = 750 / 1000 = 0.75

Government spending multiplier = 1 / (1 - MPC)

=> Government spending multiplier = 1 / (1 - 0.75) = 1 / 0.25

=> Government spending multiplier = 4

Therefore, here the government spending multiplier is 4.